OpenStudy (anonymous):

Two crews can service the space shuttle in 12 days. The faster crew can service the shuttle in 20 days alone. How long would the slower crew need to service the shuttle working alone?

OpenStudy (anonymous):

Let, X= Time taken by the slower crew, \[ \frac 1 {20} + \frac 1 x = \frac 1 {12} \implies x = ?\]

OpenStudy (anonymous):

EDIT: 'x'

OpenStudy (anonymous):

that is 30 days.